# Random Lévy Looptrees and Lévy Maps

### v2

@@ Abstract
What is the analogue of Lévy processes for random surfaces? Motivated by scaling limits of random planar maps in random geometry, we introduce and study Lévy looptrees and Lévy maps. They are defined using excursions of general Lévy processes with no negative jump and extend the known stable looptrees and stable maps, associated with stable processes. We compute in particular their fractal dimensions in terms of the upper and lower Blumenthal–Getoor exponents of the coding Lévy process. In a second part, we consider excursions of stable processes with a drift and prove that the corresponding looptrees interpolate continuously as the drift varies from $\infty$- ∞ to $\infty$, or as the self-similarity index varies from $1$ to $2$, between a circle and the Brownian tree, whereas the corresponding maps interpolate between the Brownian tree and the Brownian sphere.
@@ 1.1.   Motivation and literature
In random geometry, the Brownian sphere (also known as the Brownian map) is a random metric space almost surely homeomorphic to the two-dimensional sphere, whose study has attracted a lot of attention in the past twenty years. It has in particular been shown to be the universal scaling limit of many different models of random planar maps, see e.g. [LG13, Mie13, BLG13, BJM14, ABA17, ABA21, Abr16, CLG19, Mar22a, KM23]. We recall that a planar map is the embedding without edge-crossing of a finite, connected multigraph on the sphere, viewed up to orientation-preserving homeomorphisms. In two recent breakthroughs, it was also shown to appear at the scaling limit of models of random, non embedded, planar graphs [AFL23, Stu22]. The models of maps in the previous references share the common feature that no face has a size which dominates the other ones, so they vanish in the limit after applying a suitable rescaling. In this sense, the Brownian sphere is the canonical model of spherical random surfaces. It bears intimate connections with the Brownian excursion, Aldous’ Continuum Random Tree (hereafter called the ‘Brownian tree’), and Le Gall’s Brownian snake, which are also important objects in probability. It also relates to the theory of Liouville Quantum Gravity [MS20, MS21a, MS21b, MS21c]; let us refer to the expository papers [Mil18, Gwy20, GHS23, She23] for a gentle introduction to this topic.
@@
Parallel to this, some effort has been put into escaping the universality class of the Brownian sphere, motivated in part by statistical physics models on maps (see [LGM11, Section 8]). In this direction, Le Gall & Miermont [LGM11] constructed a one-parameter family of continuum models, called stable maps, which are the scaling limits (along subsequences) of Boltzmann random planar maps in regimes with large degree faces.
@@
where $\deg(f)$ is the degree of the face $f$. One can then sample a pointed map with $n$ edges proportionally to its weight (assuming that the total weight of $n$-edged maps, which is a finite set, is non-zero). This model is similar in spirit to the so-called model of simply generated trees, which can be seen under mild assumptions as size-conditioned Bienaymé–Galton–Watson trees (see e.g. the survey [Jan12] for details).
@@
In [KM23] we have recently considered planar maps conditioned both to have $n$ edges and $K_{n}$ vertices (and thus $n-K_{n}+2$ faces by Euler’s formula). One motivation originated from predictions in [FG14] concerning the typical order of distances in such uniform random maps (which we have confirmed by showing a stronger scaling limit result), and also by connections with random hyperbolic geometry mentioned in subsequent work [CKM22]. Another motivation was a large or moderate deviation question. Indeed, it can been shown that in the $\alpha$-stable map regime of [LGM11] as above, a map with $n$ edges has around $cn$ vertices for some constant $c\in(0,1)$ which depends on the weights $(q_{k})_{k}$. For example, $c=4\kappa$ for the particular weight sequence (1.1). It is then natural to study how the geometry changes when one forces the number of vertices to deviate from this typical behaviour. One of the contributions of [KM23] was to show that in this $\alpha$-stable regime, multiconditioning enables new continuum maps to appear by finely tuning $K_{n}$. To be concrete, in the case of (1.1), this corresponds to the regime where $n^{-1/\alpha}(K_{n}-4\kappa n)$ has a finite limit as $n\rightarrow\infty$. In this regime, one obtains subsequential scaling limits which are coded by decorated excursions of $\alpha$-stable Lévy processes with no negative jumps with a drift, see Section 6.1 for a more precise statement and Figure 1 for simulations.
@@
The objective of this present work is then twofold. First we aim at defining more general continuum maps related to unit duration excursions of any Lévy process with no negative jump. Indeed, as we have previously mentioned, the case of stable processes with a drift appears in a multiconditioned setting. More generally, we believe that the Lévy maps that we introduce and construct here are the only possible scaling limits of Boltzmann maps, based on the fact that Lévy processes are the only possible scaling limits of triangular arrays of random walks [Kal02, Theorem 16.14], that the continuous-state branching processes are the only possible scaling limits of discrete-time Bienaymé–Galton–Watson branching processes [Lam67], and that the Lévy trees are the only possible scaling limits of Bienaymé–Galton–Watson trees [DLG02]. We leave such questions for a future work, and here we focus on the construction of these Lévy maps, using ad hoc discrete models to pass to the limit.
@@
We also express the fractal dimensions of Lévy maps (Hausdorff, Minkowski, and packing dimensions) in terms of natural quantities in the underlying Lévy process, namely the Blumenthal & Getoor exponents, extending results of Le Gall & Miermont [LGM11] for $\alpha$-stable maps. Intuitively, these dimensions quantify the ‘roughness’ of these random metric spaces, and their identification is an important question in random geometry. Fractal dimensions of stable trees and stable maps have been computed respectively in [HM04, LGM11], and dimensions of general Lévy trees have been computed in [DLG05]. The key difference with Le Gall & Miermont’s approach is that self-similarity is not available anymore, and the fact that the lower and upper Blumenthal-Getoor indices may differ significantly complicates the analysis.
Interpolating between the Brownian sphere and the Brownian CRT.
Our second main objective is to study the two-parameter model of continuum maps related to stable Lévy processes with a drift that was previously mentioned. We consider limits as the index and the drift vary, and we prove that these continuum objects interpolate between the Brownian sphere and the Brownian tree.
@@ 1.2.   Main results
Throughout this work, we let $X$ be a Lévy process with no negative jump and with paths of infinite variation. We refer to [Ber96, DLG02] for details on Lévy processes and further discussions related to this work. Let us denote by $\psi\colon\lambda\in[0,\infty)\mapsto\log\mathop{}\!\mathbb{E}[\operatorname{e}^{-\lambda X_{1}}]$ its Laplace exponent; according to the Lévy–Khintchine formula, the latter takes the form:
@@
$\boldsymbol{\gamma}\coloneqq\sup\mathopen{}\mathclose{{}\left\{c\geq 0\colon\lim_{\lambda\to\infty}\lambda^{-c}\psi(\lambda)=\infty}\right\}\qquad\text{and}\qquad\boldsymbol{\eta}\coloneqq\inf\mathopen{}\mathclose{{}\left\{c\geq 0\colon\lim_{\lambda\to\infty}\lambda^{-c}\psi(\lambda)=0}\right\}.$ (1.2)
We use in this work the notation from [DLG05] and the exponents $\boldsymbol{\gamma}$ and $\boldsymbol{\eta}$ actually correspond to $\beta^{\prime}$ and $\beta$ respectively in [BG61]. Let us mention that $\boldsymbol{\eta}$ also relates to the Lévy measure $\pi$ by $\boldsymbol{\eta}=\inf\{c\geq 0\colon\int_{0}^{1}x^{c}\pi(\mathop{}\!\mathrm{d}x)<\infty\}$. We always have $\lim_{\lambda\to\infty}\lambda^{-2}\psi(\lambda)=\beta$, and since we assume that $\psi$ has paths with infinite variation, then $\lim_{\lambda\to\infty}\lambda^{-1}\psi(\lambda)=\infty$. In particular $1\leq\boldsymbol{\gamma}\leq\boldsymbol{\eta}\leq 2$; the two exponents coincide (with $\alpha$) for $\alpha$-stable processes and more generally when $\psi$ varies regularly at infinity, but they differ in general and all pairs of values $1\leq\boldsymbol{\gamma}\leq\boldsymbol{\eta}\leq 2$ are possible.
@@
Then by inverse Fourier transform $X_{t}$ admits a continuous density for every $t>0$ (even jointly continuous in time and space). These transition densities can then be used to define a regular version $X^{\mathrm{br}}$ of $(X_{t})_{t\in[0,1]}$ conditioned on $X_{0}=X_{1}=0$, which we call the bridge version of $X$. Next, by exchanging the parts prior and after the first minimum of $X^{\mathrm{br}}$, the so-called Vervaat transform allows to define an excursion $X^{\mathrm{ex}}$, which is informally a version of $(X_{t})_{t\in[0,1]}$ conditioned on $X_{0}=X_{1}=0$ and $X_{t}>0$ for every $t\in(0,1)$. Let us refer to Section 2.2 and references therein for a few details.
We can construct a looptree $\mathscr{L}(X^{\mathrm{ex}})$ from the excursion path $X^{\mathrm{ex}}$ in a way which somehow extends the construction of a tree from a continuous excursion via its contour exploration. Very informally, we see each positive jump, say $\Delta X^{\mathrm{ex}}_{t}>0$, as a vertical line segment with this length, turned into a cycle by identifying its two extremities. We then glue these cycles together by attaching the bottom $(t,X^{\mathrm{ex}}_{t-})$ of such a segment to the first point we meet on another vertical segment when going horizontally to the left, starting from $(t,X^{\mathrm{ex}}_{t-})$, at the last time $s<t$ such that $X^{\mathrm{ex}}_{s-}<X^{\mathrm{ex}}_{t-}<X^{\mathrm{ex}}_{s}$. In reality, due to the infinite variation paths, two macroscopic cycles never touch each other, and when the Gaussian parameter $\beta$ is nonzero, the looptree contains infinitesimal ‘tree parts’. In the case when $X^{\mathrm{ex}}$ is the Brownian excursion, the looptree $\mathscr{L}(X^{\mathrm{ex}})$ actually has no cycle and reduces to the usual Brownian tree. The formal definition of $\mathscr{L}(X^{\mathrm{ex}})$ is given in Section 2.2, let us only mention that it is a metric measured space, which is the quotient of the interval $[0,1]$ by a continuous pseudo-distance defined from $X^{\mathrm{ex}}$.
@@
In the discrete setting, looptrees equipped with suitable integer labels on the vertices encode pointed bipartite planar maps, see Figure 3 for an example and Section 2.1 for some details on this bijection. This corresponds to a reformulation of the celebrated bijection from [BDFG04] which is written in terms of the sort of dual two-type tree of the looptree, obtained for every cycle by joining every vertex on this cycle to an extra vertex inside. It also extends the famous Cori–Vauquelin–Schaeffer bijection for quadrangulations, in which case the looptree is simply a tree in which each edge is doubled into a cycle of length $2$.
Refer to captionRefer to captionRefer to caption
@@
We shall prove in Theorem 2.1 that when discrete excursions which encode discrete looptrees converge in distribution after suitable scaling to $X^{\mathrm{ex}}$, then the rescaled labelled looptrees converge to the looptree $\mathscr{L}(X^{\mathrm{ex}})$ equipped with a random Gaussian field of labels. This implies, by now classical arguments originally due to Le Gall [LG07], that the associated rescaled discrete maps are tight, and thus converge along subsequences to some limit metric space, see Theorem 2.4 below. We shall denote by $M$ any such subsequential limit, which we call a Lévy map associated with $X^{\mathrm{ex}}$.
@@
Almost surely, it holds
@@
Again, this extends the calculation of the dimension of the stable maps in [LGM11] as well as the Brownian sphere [LG07].
@@
The lower bounds on the dimensions follow from an upper bound on the rate of decay of the volume of balls centred at a uniform random point as the radius tends to 00. The latter is established for both looptrees and maps in Section 3. In the case of looptrees, it relies on a spinal decomposition, which rephrases the usual spinal decomposition of a tree and is formally described in terms of the coding Lévy excursion. A geometric argument, summarised in Figure 7, allows to include the ball in a simpler set defined using hitting times of Lévy processes. The core of the argument is the control of such quantities. By local absolute continuity of the excursion near an independent uniform random time with a bi-infinite Lévy process $(X_{t})_{t\in\mathbb{R}}$, we are able to transfer estimates obtained for the unconditioned process to its excursion. The volume of balls in the map is controlled similarly, with a close geometric argument, now depicted in Figure 10. The estimates are however much more challenging in the case of maps since one needs to control the Gaussian labels on the looptree.
@@
Upper bounds on the dimensions rely on Hölder continuity estimates, obtained in Section 4. Recall that both the looptree and the map are a quotient of the interval $[0,1]$ by some pseudo-distance; we prove there that the canonical projections are Hölder continuous. For looptrees, this is obtained by cutting the excursion path into small pieces, whenever it makes a jump larger than a small threshold, and controlling the variation between two such jumps; again we rely on local absolute continuity with respect to an unconditioned Lévy process. Hölder continuity estimates for the map are then easily derived from this, using the representation with labels on the looptree; by their Gaussian nature, the regularity of the labels is nearly half that of the looptree (which leads to the factor two in the dimensions).

### v3

@@ Abstract
What is the analogue of Lévy processes for random surfaces? Motivated by scaling limits of random planar maps in random geometry, we introduce and study Lévy looptrees and Lévy maps. They are defined using excursions of general Lévy processes with no negative jump and extend the known stable looptrees and stable maps, associated with stable processes. We compute in particular their fractal dimensions in terms of the upper and lower Blumenthal–Getoor exponents of the coding Lévy process. The case where the Lévy process is a stable process with a drift naturally appears in the context of stable-Boltzmann planar maps conditioned on having a fixed number of vertices and edges in a near-critical regime.
@@ 1.1.   Motivation and literature
In random geometry, the Brownian sphere (also known as the Brownian map) is a random metric space almost surely homeomorphic to the two-dimensional sphere, whose study has attracted a lot of attention in the past twenty years. It has in particular been shown to be the universal scaling limit of many different models of random planar maps, see e.g. [LG13, Mie13, BLG13, BJM14, ABA17, ABA21, Abr16, CLG19, Mar22, KM23]. We recall that a planar map is the embedding without edge-crossing of a finite, connected multigraph on the sphere, viewed up to orientation-preserving homeomorphisms. In two recent breakthroughs, it was also shown to appear at the scaling limit of models of random, non embedded, planar graphs [AFL23, Stu24]. The models of maps in the previous references share the common feature that no face has a size which dominates the other ones, so they vanish in the limit after applying a suitable rescaling. In this sense, the Brownian sphere is the canonical model of spherical random surfaces. It bears intimate connections with the Brownian excursion, Aldous’ Continuum Random Tree (hereafter called the ‘Brownian tree’), and Le Gall’s Brownian snake, which are also important objects in probability. It also relates to the theory of Liouville Quantum Gravity [MS20, MS21a, MS21b, MS21c]; let us refer to the expository papers [Mil18, Gwy20, GHS23, She23] for a gentle introduction to this topic.
@@
Parallel to this, some effort has been put into escaping the universality class of the Brownian sphere, motivated in part by statistical physics models on maps (see [LGM11, Section 8]). In this direction, Le Gall & Miermont [LGM11] constructed a family of continuum models and proved that they are scaling limits, along subsequences, of Boltzmann random planar maps (defined in the next paragraph) in regimes where the degree of the faces have infinite variance, and whose behaviour is dictated by an exponent $\alpha\in(1,2)$. Very recently, Curien, Miermont & Riera [CMR25] proved that the convergence in distribution holds without taking subsequences. They call these limits an $\alpha$-stable carpet when $\alpha\geq 3/2$ and an $\alpha$-stable gasket when $\alpha<3/2$ because of topological differences between these two regimes. To simplify, we shall call them $\alpha$-stable maps in all regimes $\alpha\in(1,2)$.
@@
where $\deg(f)$ is the degree of the face $f$. One can then sample a pointed map with $n$ edges proportionally to its weight (assuming that the total weight of $n$-edged maps, which is a finite set, is non-zero). This model is similar in spirit to the so-called model of simply generated trees, which can be seen under mild assumptions as size-conditioned Bienaymé–Galton–Watson trees, see e.g. the survey [Jan12] for details.
@@
In [KM23] we have recently considered planar maps conditioned both to have $n$ edges and $K_{n}$ vertices, and thus $n-K_{n}+2$ faces by Euler’s formula. One motivation originated from predictions in [FG14] concerning the typical order of distances in such uniform random maps, which we have confirmed by showing a stronger scaling limit result, and also by connections with random hyperbolic geometry mentioned in the subsequent work [CKM22]. Another motivation was a large or moderate deviation question. Indeed, it can be shown that in the $\alpha$-stable map regime of [LGM11] as above, a map with $n$ edges has around $\theta n$ vertices for some constant $\theta\in(0,1)$ which depends on the weights $(q_{k})_{k}$. For example, $\theta=4\kappa$ for the particular weight sequence (1.1). It is then natural to study how the geometry changes when one forces the number of vertices to deviate from this typical behaviour. One of the contributions of [KM23] was to show that in this $\alpha$-stable regime, multiconditioning enables new continuum maps to appear by finely tuning $K_{n}$. To be concrete, in the case of (1.1), this corresponds to the regime where $n^{-1/\alpha}(K_{n}-4\kappa n)$ has a finite limit as $n\rightarrow\infty$. In this regime, one obtains subsequential scaling limits which are coded by decorated excursions of $\alpha$-stable Lévy processes with no negative jumps with a drift, see Corollary 3.1 for a precise statement and Figure 1 for simulations.
@@
The objective of this present work is then twofold. First we aim at defining more general continuum maps related to unit duration excursions of any Lévy process with no negative jump. Indeed, as we have previously mentioned, the case of stable processes with a drift appears in a multiconditioned setting. More generally, we believe that the Lévy maps that we introduce and construct here are the only possible scaling limits of Boltzmann maps, based on the fact that Lévy processes are the only possible scaling limits of triangular arrays of random walks [Kal02, Theorem 16.14], that the continuous-state branching processes are the only possible scaling limits of discrete-time Bienaymé–Galton–Watson branching processes [Lam67], and that the Lévy trees are the only possible scaling limits of Bienaymé–Galton–Watson trees [DLG02]. It would also be interesting to investigate limits of gaskets of a large random $O(N)$ decorated triangulation in near critical regimes (see e.g. the discussion in [CK15, Section 5.3] which hints to the appearance of Lévy looptrees and maps). We leave such questions for a future work, and here we focus on the construction of these Lévy maps, using ad hoc discrete models to pass to the limit.
@@
We also express the fractal dimensions of Lévy maps: Hausdorff, Minkowski, and packing dimensions, in terms of natural quantities in the underlying Lévy process, namely the Blumenthal & Getoor exponents, extending results of Le Gall & Miermont [LGM11] for $\alpha$-stable maps. Intuitively, these dimensions quantify the ‘roughness’ of these random metric spaces, and their identification is an important question in random geometry. Fractal dimensions of stable trees and stable maps have been computed respectively in [HM04, LGM11], and dimensions of general Lévy trees have been computed in [DLG05]. The key difference with Le Gall & Miermont’s approach is that self-similarity is not available anymore, and the fact that the lower and upper Blumenthal-Getoor indices may differ significantly complicates the analysis.
@@ 1.2.   Main results
Throughout this work, we let $X=(X_{t})_{t\geq 0}$ be a Lévy process with no negative jump and with paths of infinite variation. We refer to [Ber96, DLG02] for details on Lévy processes and further discussions related to this work. Let us denote by $\psi\colon\lambda\in[0,\infty)\mapsto\log\mathop{}\!\mathbb{E}[\operatorname{e}^{-\lambda X_{1}}]$ its Laplace exponent; according to the Lévy–Khintchine formula, the latter takes the form:
@@
$\boldsymbol{\gamma}$ $\coloneqq\sup\Bigl{\{}c\geq 0\colon\lim_{\lambda\to\infty}\lambda^{-c}\psi(\lambda)=\infty\Bigr{\}},$ (1.2)
$\boldsymbol{\eta}$ $\coloneqq\inf\Bigl{\{}c\geq 0\colon\lim_{\lambda\to\infty}\lambda^{-c}\psi(\lambda)=0\Bigr{\}}.$
We use in this work the notation from [DLG05] and the exponents $\boldsymbol{\gamma}$ and $\boldsymbol{\eta}$ actually correspond to $\beta^{\prime\prime}$ and $\beta$ respectively in [BG61]. Let us mention that $\boldsymbol{\eta}$ also relates to the Lévy measure $\pi$ by $\boldsymbol{\eta}=\inf\{c\geq 0\colon\int_{0}^{1}x^{c}\pi(\mathop{}\!\mathrm{d}x)<\infty\}$. We always have $\lim_{\lambda\to\infty}\lambda^{-2}\psi(\lambda)=\beta$, and since we assume that $X$ has paths with infinite variation, then $\lim_{\lambda\to\infty}\lambda^{-1}\psi(\lambda)=\infty$. In particular $1\leq\boldsymbol{\gamma}\leq\boldsymbol{\eta}\leq 2$; the two exponents coincide (with $\alpha$) for $\alpha$-stable processes and more generally when $\psi$ varies regularly at infinity, but they differ in general and all pairs of values $1\leq\boldsymbol{\gamma}\leq\boldsymbol{\eta}\leq 2$ are possible.
@@
As observed by Kallenberg [Kal81, Section 5], this condition holds as soon as either $\beta>0$ or $u^{-2}|\log u|^{-1}\int_{0}^{u}r^{2}\pi(\mathop{}\!\mathrm{d}r)\to\infty$ as $u\to 0$. Observe that this last convergence requires only slightly more than $\pi$ to be infinite: $\pi$ is infinite as soon as $u^{-2}\int_{0}^{u}r^{2}\pi(\mathop{}\!\mathrm{d}r)\to\infty$, and is finite as soon as $u^{-2}|\log u|^{1+\varepsilon}\int_{0}^{u}r^{2}\pi(\mathop{}\!\mathrm{d}r)\to 0$.
Condition (1.3) also appears in [Kni96, UB14], and enables us to define bridges and excursions of the process with a fixed duration. Precisely, by inverse Fourier transform, under (1.3), the random variable $X_{t}$ admits a continuous density for every $t>0$ (even jointly continuous in time and space). These transition densities can then be used to define a regular version $X^{\mathrm{br}}$ of $(X_{t})_{t\in[0,1]}$ conditioned on $X_{0}=X_{1}=0$, which we call the bridge version of $X$. Next, by exchanging the parts prior and after the first minimum of $X^{\mathrm{br}}$, the so-called Vervaat transform allows to define an excursion $X^{\mathrm{ex}}$, which is informally a version of $(X_{t})_{t\in[0,1]}$ conditioned on $X_{0}=X_{1}=0$ and $X_{t}>0$ for every $t\in(0,1)$. Let us refer to Section 2.2 and references therein for a few details. Without the condition (1.3), the construction of the continuum looptrees and maps as well as the calculation of their dimensions can be carried out in the setting of unconditioned Lévy processes, or under the Itō excursion measure.
We can construct a looptree $\mathscr{L}(X^{\mathrm{ex}})$ from the excursion path $X^{\mathrm{ex}}$ in a way that somehow extends the construction of a tree from a continuous excursion via its contour exploration. Very informally, we see each positive jump, say $\Delta X^{\mathrm{ex}}_{t}>0$, as a vertical line segment with this length, turned into a cycle by identifying its two extremities. We then glue these cycles together by attaching the bottom $(t,X^{\mathrm{ex}}_{t-})$ of such a segment to the first point we meet on another vertical segment when going horizontally to the left, starting from $(t,X^{\mathrm{ex}}_{t-})$, at the last time $s<t$ such that $X^{\mathrm{ex}}_{s-}<X^{\mathrm{ex}}_{t-}<X^{\mathrm{ex}}_{s}$. In reality, due to the infinite variation paths, two macroscopic cycles never touch each other, and when the Gaussian parameter $\beta$ is nonzero, the looptree contains infinitesimal ‘tree parts’. In the case when $X^{\mathrm{ex}}$ is the Brownian excursion, the looptree $\mathscr{L}(X^{\mathrm{ex}})$ actually has no cycle and reduces to a scaled version of the usual Brownian tree. The formal definition of $\mathscr{L}(X^{\mathrm{ex}})$ is given in Section 2.2, let us only mention that it is a metric measured space, which is the quotient of the interval $[0,1]$ by a continuous pseudo-distance defined from $X^{\mathrm{ex}}$.
@@
Turning to Lévy maps, one could construct them directly from the process $X^{\mathrm{ex}}$ together with extra randomness given by a Gaussian field on the looptree $\mathscr{L}(X^{\mathrm{ex}})$, using the analogue of the formula of $D^{\ast}$ in [LG07, LG13, Mie13] to define the Brownian sphere. However this is not what we will do and, as for the pioneer works in the Brownian and stable regimes [LG07, LGM11], we will instead use approximations by rescaled discrete maps. The reason is twofold: first our motivation to study these objects in the first place is to eventually prove scaling limit results, and second this will allow us to use the discrete objects to obtain some estimates that seem harder to obtain directly in the continuum world.
We shall give more details in Section 2 below, but for the benefit of the reader let us here give a first rather informal sense of these Lévy maps. A celebrated bijection from [BDFG04] shows that discrete planar maps with a distinguished vertex are encoded by certain trees whose vertices are equipped with suitable integer labels. We can reformulate this bijection in terms of looptrees, again with labelled vertices, as shown in Figure 3. Then just as in the continuous setting, a discrete looptree can be coded by a discrete excursion $W^{n}$ of duration $n$ (the number of edges of the maps) with increments in $\mathbb{Z}_{\geq-1}=\{-1,0,1,2,\dots\}$. Under mild exchangeability conditions, we shall prove in Theorem 2.1 that when such random excursions converge in distribution after suitable scaling to $X^{\mathrm{ex}}$, then the rescaled looptrees converge to the looptree $\mathscr{L}(X^{\mathrm{ex}})$ for the so-called Gromov–Hausdorff–Prokhorov topology. In addition, the random labels on the vertices of the looptree converge in the scaling limit to a random Gaussian field on $\mathscr{L}(X^{\mathrm{ex}})$.
@@
By now classical arguments originally due to Le Gall [LG07], this implies that the discrete random maps associated with these labelled looptrees are tight, and thus converge along subsequences to some limit metric space. Precisely, we shall prove in Theorem 2.3 below a formal version of the following statement.
‘Definition’ of the Lévy maps, to be made more precise in Theorem 2.3.
Let $M^{n}$ be random bipartite planar map with $n$ edges and a distinguished vertex and let $W^{n}$ denote the discrete excursion that codes its associated looptree. Assume an exchangeability property for the increments of $W^{n}$ (i.e. for the faces of $M^{n}$), and that there exists a sequence $r_{n}\to\infty$ such that the rescaled path $(r_{n}^{-1}W^{n}_{\lfloor nt\rfloor})_{t\in[0,1]}$ converges in distribution to $X^{\mathrm{ex}}$ for the Skorokhod topology. Then from every increasing sequence of integers, one can extract a subsequence along which the rescaled maps $(2r_{n})^{-1/2}M^{n}$ converge in distribution to some non degenerate measured metric space for the Gromov–Hausdorff–Prokhorov topology. We shall denote by $M$ any such subsequential limit, which we call a Lévy map associated with $X^{\mathrm{ex}}$.
Theorem 2.3 actually provides more information and shows for example that the graph distances in $M^{n}$ to its distinguished vertex converge in distribution, without the need to extract a subsequence. However the convergence of all the pairwise distances remains open. This result is the generalisation of [LG07, LGM11] in the Brownian and stable regimes, before one was able to prove that convergence of maps holds without the need to extract a subsequence. These works also compute the Hausdorff dimensions of the subsequential limits and our next theorem extends this to the Lévy regime.
@@
Almost surely, for any subsequential limit $M$, it holds
@@
The lower bounds on the dimensions follow from an upper bound on the rate of decay of the volume of balls centred at a uniform random point as the radius tends to 00. The latter is established for both looptrees and maps in Section 4. In the case of looptrees, it relies on a spinal decomposition, which rephrases the usual spinal decomposition of a tree and is formally described in terms of the coding Lévy excursion. A geometric argument, summarised in Figure 6, allows to include the ball in a simpler set defined using hitting times of Lévy processes. The core of the argument is the control of such quantities. By local absolute continuity of the excursion near an independent uniform random time with a bi-infinite Lévy process $(X_{t})_{t\in\mathbb{R}}$, we are able to transfer estimates for that set obtained for the unconditioned process to its excursion. The volume of balls in the map is controlled similarly, with a close geometric argument, now depicted in Figure 9. The estimates however are much more challenging in the case of maps, since here one needs to control the Gaussian labels on the looptree.
@@
Upper bounds on the dimensions rely on Hölder continuity estimates, obtained in Section 5. Indeed both the looptree and the map are a quotient of the interval $[0,1]$ by some pseudo-distance, and we prove there that the canonical projections are Hölder continuous. For looptrees, this is obtained by cutting the excursion path into small pieces, whenever it makes a jump larger than a small threshold, and controlling the variation between two such jumps; again we rely on local absolute continuity with respect to an unconditioned Lévy process. Hölder continuity estimates for the map are then easily derived from this, using the representation with labels on the looptree; by their Gaussian nature, the regularity of the labels is nearly half that of the looptree, which leads to the factor two in the dimensions.